I need the bottom part doneOn January 1, 2021, Whispering Co. issued ten-year bonds with a face value of $4,900,000 and a stated interest rate of 10%, p
Answer #1

Face value = $4,900,000

Semiannual interest : $4,900,000*10%*6/12 = $245,000

Market Rate = 12%, Semiannual = 6%

semi annual Payments = 10*2 =20

Issue Price of the Bond = 4,900,000*0.312+245,000*11.470 = $4,338,950

Issue Price = 4,338,950

Cash interest =4900000*10%*6/12 245000 Interest Expenses4331600*12%*6/12 259896 4346496*125*6/12 260790 Amortization Sechedul